OpenStudy (anonymous):

What are the zero(s) of the function f(x) = 5x^2 - 25x / x

OpenStudy (solomonzelman):

\[5x^2 - 25x / x\]is same as\[5x^2-25\]\[5x^2-25=0\]\[5(x^2-5)=0\]\[x^2-5=0\]\[x^2=5\]\[x=±\sqrt{5}\]

OpenStudy (solomonzelman):

both of them work, and both are not real roots.

OpenStudy (anonymous):

It is 0 and 5? @SolomonZelman what happened between 5(x^2−5)=0 and x^2−5=0? Where did the 5 go?

OpenStudy (solomonzelman):

I divide both sides by 5.

OpenStudy (solomonzelman):

\[5(x^2-5)\color{red} {\div 5}=0\color{red} {\div 5}\]\[x^2-5=0 \color{red} {\div 5}=0\]

OpenStudy (anonymous):

Ohh, okay. Sorry, I should have known that. Thank you for your help! @SolomonZelman

OpenStudy (solomonzelman):

Anytime!